OpenStudy (anonymous):

they is a sale of 20% and the price now is 32 what is the original price ?

OpenStudy (anonymous):

Given: sale = 20% After discount, price = 32 RTF ( required to find ) original price. To find the discounted price, we multiply the original price with the sale , but we don't know the original price so we'll call it "x" in here ^_^ : (original Price)(20%) = (discounted Price) \[x \frac{20}{100} = 32 \rightarrow 20x =(32)(100)\] \[x = \frac{(100)(32)}{20} \rightarrow x = 160\] so the original price was $160 is it clear now? :)